### v1

@@ 3 Department of Astronomy, Westlake University, Hangzhou, Zhejiang 310030, China
We introduce a novel dynamical model, named empirical triaxial orbit-superposition model, for the Milky Way halo. This model relies on minimal physical assumptions that the system is stationary, meaning the distribution function in 6D phase-space does not change when the stars orbiting in the correct gravitational potential. We validate our method by applying it to mock datasets that mimic the observations of the Milky Way halo from LAMOST + Gaia with stars’ 3D position and 3D velocity observed. By removing the stellar disk and substructures, correcting the selection function, we obtain a sample of smooth halo stars considered as stationary and complete. We construct a gravitational potential including a highly flexible triaxial dark matter halo with adaptable parameters. Within each specified gravitational potential, we integrate orbits of these halo stars, and build a model by superposing the orbits together taking the weights of stars derived from the selection function correction. The goodness of the models are evaluated by comparing the density distributions as well as 3D velocity distributions numerically represented in the model to that in the data. The shape and radial density distribution of the underlying dark matter halo can be constrained well simultaneously. We apply it to three mock galaxies with different intrinsic shapes of their dark matter halos and achieved accurate recovery of the 3D dark matter density distributions for all.
@@ 1 Introduction
The three-dimensional (3D) distribution of dark matter (DM) within the Milky Way (MW) is crucial for understanding its formation history. Despite extensive studies, there remain significant uncertainties regarding its total mass, density profile, and particularly its debated 3D shape.
@@
Streams, viewed as simple orbital pathways that trace the halo, played a significant role in constraining the shape of the MW DM halo. Several approaches have aimed to constrain the DM flattening using the two cold streams, Pal 5 and GD–1, which generally suggest a consistent result of $q=0.8-0.95$, assuming an axisymmetric halo (Bovy et al., 2016; Bowden et al., 2015; Koposov et al., 2010; Küpper et al., 2015; Malhan & Ibata, 2019). These streams trace the inner halo at $r\lesssim$ 20 kpc. However, some other streams in similar regions indicate a preference for a more prolate halo. Recent research involving the Heimi stream has identified a mildly triaxial dark matter halo, with parameters $p=1.013$ and $q=1.204$ within the inner 20 kpc (Woudenberg & Helmi, 2024). Another stream, M68 (Palau & Miralda-Escudé, 2019), might also favor a prolate halo. The combination of data from NGC 3201, M68, and Palomar 5 fits well with a halo axis ratio $q=1.06$. (Palau & Miralda-Escudé, 2023).
@@
In the outer halo, the Sagittarius stream, being the most notable structure, is analyzed using a variety of methods, which results in different outcomes with $q$ ranges from 0.44 to 1.3 Panithanpaisal et al. (2022); Law & Majewski (2010); Johnston et al. (2005); Helmi (2004); Ibata et al. (2001). These inconsistent results could be mostly due to different parametrizations of the potential and the treatment of the influence of the LMC on the outer portion of the stream (Vera-Ciro & Helmi, 2013). Recent research shows that when modeling the Sagittarius stream in a dynamic potential, acknowledging our Galaxy’s motion towards the massive in-falling LMC, the shape of the dark matter halo is not strongly constrained by the Sagittarius stream (Vasiliev et al., 2021).
@@
Conversely, halo stars–with full 6D phase-space information observed (3D position and 3D velocities)–span the entire space and, statistically, should exhibit stronger power in constraining the underlying DM distribution. Nevertheless, these stars exist with complex orbital dynamics, necessitating accurate dynamical modeling. The current dynamical models applied to the MW halo usually impose ad hoc assumptions that might bias the results.
@@
A commonly used dynamical model for the MW halo is the Jeans model, which often assumes either a spherical or axisymmetric configuration. Spherical Jeans models, which utilize giant stars in the halo extending to large radii $r\sim 100$ kpc, provide strong constraints on the MW halo mass (e.g., Xue et al., 2008; Bird et al., 2022), but do not restrict its shape by definition. Axisymmetric assumptions in the Jeans model generally yield mass estimates for the MW that are consistent with those from the spherical model (see review Wang et al. (2020)). However, the results diverge regarding the DM halo’s shape. Using RR Lyrae stars, research by Loebman et al. (2014) identified an oblate halo with $q=0.4\pm 0.1$ at $r<20$ kpc, but Wegg et al. (2019) found a nearly spherical halo with $q=1.00\pm 0.09$ at $205<r<$ 20 kpc. The traditional axisymmetric Jeans model’s alignment of the velocity ellipsoid differs from observations of MW halo stars, potentially leading to bias. Recently, an axisymmetric Jeans model that allows for a generic shape and radial orientation of the velocity ellipsoid has been developed (Nitschai et al., 2020) and applied to MW BHB and K-giant stars, resulting in a near spherical DM halo with $q=0.9-0.95$ (Zhang et al., 2025). Nevertheless, Jeans models are still limited in axisymmetric assumption and that they do not use the full 6D phase-space information, as they often only fit the first and second velocity moments and assume a specific tracer density.
@@
To effectively use the 6D phase space data, distribution functions (DF) based dynamical models have been developed and implemented for the MW. A DF-based dynamic model applied to globular clusters (GCs) by (Posti & Helmi, 2019) identifies a prolate halo with $q=1.30\pm 0.25$. Conversely, using DF-based axisymmetric models on halo stars reveals an oblate halo with $q>0.963$ within $r<30$ kpc, as determined by RR Lyrae stars (Hattori et al., 2021; Li & Binney, 2022). These models presume an analytically assumed DF. A new approach named the Empirical Distribution Function has been proposed recently; it uses the observed 6D data of the tracers to construct the DF through appropriate smoothing of the observational data based on theoretical insights, making the DF data-driven with fewer inherent assumptions. Nevertheless, this model is restricted to spherical symmetry (Li et al., 2024). Recently, a novel method was developed for determining the gravitational potential of the MW using 6D data, making use of tools of deep learning (Green et al., 2023), taking minimal physical assumptions that the system is being stationary. This method learns the DF from the data and does not restrict on the gravitational potential; however, it is not applied to the MW halo yet as it may require ×$\times 10^{5}$ data points to construct the model.
@@
Currently, there isn’t a triaxial dynamical model that properly represents the MW halo. However, triaxial particle-based or orbit-based models are widely used for nearby galaxies and the MW bulge. In such models, the DF of the tracer is numerically constructed by superposition of stellar orbits. For nearby galaxies, the observational information is integrated along the line-of-sight, without resolution of single stars; the true stellar orbit distribution is unknown. In the traditional Schwarzschild orbit-superposition method, a comprehensive orbit library is created based on theoretical assumptions, and orbit weights are determined by fitting to the observational data, i.e., line-of-sight velocity distribution across the 2D projected plane (e.g., van den Bosch et al., 2008; Zhu et al., 2018; Thater et al., 2022; Tahmasebzadeh et al., 2024; Vasiliev & Athanassoula, 2015). The limited observational data restricts the precision of the DF composed of stellar orbits, and the ability to determine the DM shape of nearby galaxies. In the MW, we have measurements of the 6D phase space for individual stars. However, the bulge areas are significantly affected by dust extinction, leading to incomplete observations. Analytical models continue to be employed to represent the stellar density distribution and used as model input. Libraries of orbits based on theoretical assumptions, alongside the determination of orbit weights through data fitting, remain necessary for modeling the MW bulge (Khoperskov et al., 2024a; Portail et al., 2017; Long et al., 2013; Khoperskov et al., 2024c) and disk (Khoperskov et al., 2024b; Bovy et al., 2018).
@@
In this paper, we introduce an empirical triaxial orbit-superposition model for the Milky Way halo with the orbit library and orbit weights entirely determined from the observations. We thus take the minimum physical assumption that the stellar halo is stationary, meaning the DF of stars integrated in a correct potential matches that of stars directly observed instantaneously. We apply the method to mock datasets mimicking the LAMOST and Gaia observations. Constrained by halo stars in 6D phase space, we will show that the 3D shape and radial density distribution of the Milky Way’s DM halo can be rigorously determined simultaneously, and the results are highly data-driven.
The paper is organized in the following way: we describe the details of mock data created from Auriga simulations in Section 2, we introduce the method and its application to the mock data in Section 3, we show the results in Section 4 and discuss them in Section 5. We conclude in Section 6.
@@ 2 Simulations and mock data
We first create mock data mimicking the LAMOST and Gaia observations from Auriga simulations. Then we process and refine the data to ready it for dynamical modeling, following the same procedures applied to real observations.
@@ 2.1.1 Coordinates and measurement of true DM shape
The DM halo typically does not align precisely with the stars. The principal axes of the DM halo, denoted as $\boldsymbol{X}\equiv\{X,Y,Z\}$, may be connected to the galactocentric coordinate through a rotation matrix, characterized by three Euler angles: $\alpha_{q}$, $\beta_{q}$, and $\gamma_{q}$:
@@
In practice, when deriving DM distributions from observational data, the DM orientation angles and intrinsic shapes are highly degenerate. To streamline this, we set $Z=z_{\rm gc}$ by assuming $\beta_{q}=0$. The DM halo’s orientation in the $x_{\rm gc}y_{\rm gc}$ plane is left unrestricted. By projecting DM particles along the $Z$ axis, we determine the principal axis in the $XY$ plane using particles within $r<100$ kpc. The long axis in this plane is labeled as $X$ and the short as $Y$. We define the axis ratio as $p_{\rm DM}=Y/X$, and $q_{\rm DM}=Z/X$. In this definition, $p_{\rm DM}$ is restricted to be $<1$, while $q_{\rm DM}$ is allowed to be $>1$.
The intrinsic shape of Auriga halo 23, 5, and 12 measured in such a coordinate $XYZ$ is shown in Fig. 1. The intrinsic shapes measured in our coordinate $XYZ$ are slightly different but close to the intrinsic shapes that are measured in Prada et al. (2019).
@@ 2.2 Mock data
By cross-matching LAMOST with Gaia DR3, we have a sample of 619,284 k-giants with 3D velocity measured, the typical errors of radial velocity $v_{r}$, tangential velocity $v_{\phi}$, $v_{\theta}$ are 10, 30, 30 km/s, respectively. The typical error of distance is about $15\%$. We exclude disk stars chemo-dynamically to avoid biasing the spatial distribution of the halo. We carefully exclude all substructures, including streams, overdensities, and GES using the friend-of-friend method (Wang et al., 2022; Yang et al., 2019), and further perform $3\sigma$ clipping to remove the outliers. At the end, we remain with 15,000 k-giants in the smooth halo with spatial distribution shown in Fig. 2.
@@
First, we treat each particle as a single star, transferring their positions and velocities into the ICRS coordinate. We adopt a relative error of $15\%$ for distance, an error of 30 km/s for the tangential velocity in RA and DEC directions, and an error of 20 km/s for the LOS velocity. The positions and velocities of stars are then perturbed by random numbers, normally distributed with dispersions equal to the observational errors.
Second, we perform spatial selection of the stars. We divide the particles into an original bulge with $r_{\rm gc}<4$ kpc, an original disk with $z_{\rm gc}<4$ kpc, $r_{\rm gc}>4$ kpc and $Z/Z_{\odot}>0.5$, and an original halo with the rest. We first randomly select 1/100 particles from the disk, 1/3 particles from the halo (1/4 for Auriga 12), and no particles from the bulge, then perform further spatial selection by only including stars in the northern hemisphere and within a spatial volume similar to the MW LAMOST + Gaia observations, as shown in Fig.2. This results in $\sim 30000$ stars with full 6D phase space information measured. The sample size is much smaller than the original LAMOST + Gaia sample because we take sparse sampling of the disk stars.
@@ 2.3 Data preparation and correction of selection function
With mock data created in the observational manner, we convert the stars into the galactocentric coordinate and obtain the uncertainty of position and velocity in the new coordinate by a bootstrapping process. We clean the data and correct the select function to make it prepared for the model construction, in a way similar to how we deal with the real observations.
@@
First, we use a position-velocity clustering estimator method to identify substructures; we select the smooth halo by removing all the stars identified in substructures with a linking length of 0.84 in the 6D distance (Yang et al., 2019). The stellar halos of Auriga 5 and 23 are rather smooth with no significant structures; we only perform this step on the mock data of Auriga 12. There are fewer substructures in the halo of Auriga simulations than in the real Milky Way galaxy due to the limited mass resolution in the simulation.
Second, we remove the stars from the disk. We integrate the stellar orbits in an initially guessed potential and calculate the circularity $\lambda_{z}$ of each orbit. $\lambda_{z}$ is defined as the angular momentum around the z axis normalized by the maximum that is allowed by circular orbits with the same binding energy:$\lambda_{z}=\overline{L_{z}}/(rV_{c})$, where $\overline{L_{z}}=\overline{xv_{y}-yv_{x}}$, $r=\sqrt{\overline{x^{2}+y^{2}+z^{2}}}$, and $V_{c}=\sqrt{\overline{v_{x}^{2}+v_{y}^{2}+v_{z}^{2}}}$, taking the average of the points along each orbit. A star on a near circular orbit has $\lambda_{z}\sim 1$, and on a radial orbit has $\lambda_{z}\sim 0$. We remove the disk stars in a cut combining circularity and metallicity; all the stars with either $\lambda_{z}>0.8$ or $\lambda_{z}>0.5$ and $Z/Z_{\odot}>0.5$ are removed. As shown in Fig. 3, the remaining halo stars have almost a symmetric distribution in $\lambda_{z}$ centered at $\lambda_{z}=0$.
@@
Third, we clean the data by $3\sigma$ clipping in the velocity distributions. Although we have removed the substructures identified by the clustering estimator method, there are still some particles with high tangential velocities which are unlikely to be part of a smooth halo in dynamical equilibrium. We further perform a $3\sigma$ clipping in $v_{\phi}$ versus $v_{\theta}$ to remove the outliers.
@@
Lastly, we correct the selection function. For this proof-of-concept, we take a simple approach by correcting the selection function in the $R_{\rm gc}-z_{\rm gc}$ plane. We divide the $R_{\rm gc}-z_{\rm gc}$ plane into $50\times 50$ bins in $R_{\rm gc}=[0,50]$ kpc and $z_{\rm gc}=[0,50]$ kpc. We calculate the density of stars in each volume of $2\pi R_{\rm gc}dR_{\rm gc}dz_{\rm gc}$ constructed from our sample ($\rho_{\rm select}$) and that from the original halo before performing spatial selection ($\rho_{\rm original}$). Any star $i$ in the bin $j$ will obtain a weight of $w_{i}=\rho_{\rm original,j}/\rho_{\rm select,j}$. Such correction of selection is less accurate at $z_{\rm gc}<4$ kpc where disk stars are mixed in the original halo.
@@
At the end, we have about 20,000 stars in the smooth halo; these stars are all located in the northern hemisphere and have the selection function corrected. The final sample of stars with their weights is shown in Fig 2.
@@
The real selection function of LAMOST is complicated, and it is corrected by comparing to photometric data which are considered complete with magnitude limit (Liu et al., 2017; Yang et al., 2022), which results in the correction of the selection function in the 3D space, more accurate than what we did here. It is beyond the scope of this paper to test the method of selection function correction. We just wanted the spatial sampling of our mock data to match some major properties of the real observations: (1) the data is not complete in different azimuthal angles, and we have less data in the direction of the Galactic Center; (2) the observations at larger distances or with small polar angles are less complete; thus, we got higher weights for the stars at these positions. As shown in Fig. 2, the particle weight distribution in our mock dataset generally matches the trend in the LAMOST + Gaia data of the MW.

### v2

@@ 3 Department of Astronomy, Westlake University, Hangzhou, Zhejiang 310030, China
We introduce a novel dynamical model, named empirical triaxial orbit-superposition model, for the Milky Way halo. This model relies on minimal physical assumptions that the system is stationary, meaning the distribution function in 6D phase-space does not change when the stars orbit in the correct gravitational potential. We validate our method by applying it to mock datasets that mimic the observations of the Milky Way halo from LAMOST + Gaia with stars’ 3D position and 3D velocity observed. By removing the stellar disk and substructures and correcting the selection function, we obtain a sample of smooth halo stars considered stationary and complete. We construct a gravitational potential including a highly flexible triaxial dark matter halo with adaptable parameters. Within each specified gravitational potential, we integrate orbits of these halo stars and build a model by superposing the orbits together taking the weights of stars derived from the selection function correction. The goodness of the models are evaluated by comparing the density distributions as well as 3D velocity distributions numerically represented in the model to that in the data. The shape and radial density distribution of the underlying dark matter halo can be constrained well simultaneously. We applied it to three mock galaxies with different intrinsic shapes of their dark matter halos and achieved accurate recovery of the 3D dark matter density distributions for all.
@@ 1 Introduction
The three-dimensional (3D) distribution of dark matter (DM) within the Milky Way (MW) is crucial to understanding its formation history. Despite extensive studies, there remain significant uncertainties regarding its total mass, density profile, and, in particular, its debated 3D shape.
@@
Streams, viewed as simple orbital pathways that trace the halo, played a significant role in constraining the shape of the MW DM halo. Several approaches have aimed to constrain the flattening of the DM using the two cold streams, Pal 5 and GD–1, which generally suggest a consistent result of q=0.80.95q=0.8-0.95, assuming an axisymmetric halo (Bovy et al., 2016; Bowden et al., 2015; Koposov et al., 2010; Küpper et al., 2015; Malhan & Ibata, 2019). These streams trace the inner halo at $r\lesssim$ 20 kpc. However, some other streams in similar regions indicate a preference for a more prolate halo. Recent research involving the Heimi stream has identified a mildly triaxial dark matter halo, with parameters $p=1.013$ and $q=1.204$ within the inner 20 kpc (Woudenberg & Helmi, 2024). Another stream, M68 (Palau & Miralda-Escudé, 2019), might also favour a prolate halo. The combination of data from NGC 3201, M68, and Palomar 5 fits well with a halo axis ratio $q=1.06$. (Palau & Miralda-Escudé, 2023).
@@
In the outer halo, the Sagittarius stream, being the most notable structure, is analysed using a variety of methods, which results in different outcomes with $q$ ranges from 0.44 to 1.3 Panithanpaisal et al. (2022); Law & Majewski (2010); Johnston et al. (2005); Helmi (2004); Ibata et al. (2001). These inconsistent results could be mainly due to different parameterizations of the potential and treatment of the influence of the LMC on the outer portion of the stream (Vera-Ciro & Helmi, 2013). Recent research shows that when modelling the Sagittarius stream in a dynamic potential, acknowledging our Galaxy’s motion towards the massive in-falling LMC, the shape of the dark matter halo is not strongly constrained by the Sagittarius stream (Vasiliev et al., 2021).
@@
Conversely, halo stars–with full 6D phase-space information observed (3D position and 3D velocities)–span the entire space and, statistically, should exhibit stronger power in constraining the underlying DM distribution. However, these stars exist with complex orbital dynamics, necessitating accurate dynamical modeling. The current dynamical models applied to the MW halo usually impose ad hoc assumptions that might bias the results.
@@
A commonly used dynamical model for the MW halo is the Jeans model, which often assumes either a spherical or axisymmetric configuration. Spherical Jeans models, which utilise giant stars in the halo extending to large radii $r\sim 100$ kpc, provide strong constraints on the MW halo mass (e.g., Xue et al., 2008; Bird et al., 2022), but do not restrict its shape by definition. The axisymmetric assumptions in the Jeans model generally yield mass estimates for the MW that are consistent with those of the spherical model (see review Wang et al. (2020)). However, the results diverge with respect to the DM halo’s shape. Using RR Lyrae stars, research by Loebman et al. (2014) identified an oblate halo with $q=0.4\pm 0.1$ at $r<20$ kpc, but Wegg et al. (2019) found a nearly spherical halo with $q=1.00\pm 0.09$ at $5<r<20$ kpc. The alignment of the velocity ellipsoid in the traditional axisymmetric Jeans model differs from observations of MW halo stars, potentially leading to bias. Recently, an axisymmetric Jeans model that allows for a generic shape and radial orientation of the velocity ellipsoid has been developed (Nitschai et al., 2020) and applied to MW BHB and K-giant stars, resulting in a near spherical DM halo with q=0.90.95q=0.9-0.95 (Zhang et al., 2025). Nevertheless, Jeans models are still limited in axisymmetric assumption and do not use the full 6D phase-space information, as they often only fit the first and second velocity moments and assume a specific tracer density.
@@
To effectively use the 6D phase space data, distribution functions (DF) based dynamical models have been developed and implemented for the MW. A DF-based dynamical model applied to globular clusters (GCs) by (Posti & Helmi, 2019) identifies a prolate halo with $q=1.30\pm 0.25$. In contrast, using DF-based axisymmetric models on halo stars reveals an oblate halo with $q>0.963$ within $r<30$ kpc, as determined by RR Lyrae stars (Hattori et al., 2021; Li & Binney, 2022). These models presume an analytically assumed DF. A new approach named the Empirical Distribution Function has been proposed recently; it uses the observed 6D data of the tracers to construct the DF through appropriate smoothing of the observational data based on theoretical insights, making the DF data-driven with fewer inherent assumptions. Nevertheless, this model is restricted to spherical symmetry (Han et al., 2016; Li et al., 2024). Recently, a novel method was developed to determine the gravitational potential of the MW using 6D data, using deep learning tools (Green et al., 2023), making minimal physical assumptions that the system is stationary. This method learns the DF from the data and does not restrict the gravitational potential; however, it has not been applied to the MW halo yet, as it may require $\times 10^{5}$ data points to construct the model.
@@
Currently, there is no triaxial dynamical model that properly represents the MW halo. However, triaxial particle-based Made-to-Measure (Portail et al., 2017; Long et al., 2013) and orbit-based Schwarzschild models (Khoperskov et al., 2024a) are widely used for nearby galaxies and the MW bulge. In such models, the DF of the tracer is numerically constructed by superposition of stellar orbits. For nearby galaxies, the observational information is integrated along the line-of-sight, without resolution of single stars; the true stellar orbit distribution is unknown. In the traditional Schwarzschild orbit-superposition method, a comprehensive orbit library is created based on theoretical assumptions, and orbit weights are determined by fitting to the observational data, i.e., line-of-sight velocity distribution across the 2D projected plane (e.g., van den Bosch et al., 2008; Zhu et al., 2018; Thater et al., 2022; Tahmasebzadeh et al., 2024; Vasiliev & Athanassoula, 2015). The limited observational data restrict the precision of the DF composed of stellar orbits and the ability to determine the DM shape of nearby galaxies. In the MW, we have measurements of the 6D phase space for individual stars. However, the bulge areas are significantly affected by dust extinction, leading to incomplete observations. Analytical models continue to be employed to represent the stellar density distribution and used as model input. Libraries of orbits based on theoretical assumptions, along with the determination of orbit weights through data fitting, remain necessary to model the MW bulge (Khoperskov et al., 2024a; Portail et al., 2017; Long et al., 2013; Khoperskov et al., 2024c) and disk (Khoperskov et al., 2024b; Bovy et al., 2018).
@@
In this paper, we introduce an empirical triaxial orbit-superposition model for the Milky Way halo. Comparing to traditional Made-to-Measure and Schwarzschild method, we do not theoretically sample the orbit/particle libraries or determine the particle/orbit weights by fitting the model to data. Instead, we integrate the orbits of stars with 6D phase-space information observed and take the weights of stars from correction of selection function, i.e., the orbit library and orbit weights are entirely determined from the observations. We take the minimum physical assumption that the stellar halo is stationary, meaning the DF of stars integrated in a correct potential matches that of stars directly observed instantaneously, our model results will thus be highly data driven. However, this method requires high-quality observations and is confined to the Milky Way only.
We apply the method to mock datasets that mimic the LAMOST and Gaia observations. Constrained by halo stars in 6D phase space, we show that the 3D shape and the radial density distribution of the Milky Way DM halo can be determined simultaneously. The paper is organised in the following way: we describe the details of mock data created from Auriga simulations in Section 2, we introduce the method and its application to the mock data in Section 3, we show the results in Section 4 and discuss them in Section 5. We conclude in Section 6.
@@ 2 Simulations and mock data
We first create mock data mimicking the LAMOST and Gaia observations from Auriga simulations. Then we process and refine the data to prepare them for dynamical modelling, following the same procedures applied to real observations.
@@ 2.1.1 Coordinates and measurement of true DM shape
The DM halo typically does not align precisely with the stars. The principal axes of the DM halo, denoted $\boldsymbol{X}\equiv\{X,Y,Z\}$, may be connected to the galactocentric coordinate through a rotation matrix, characterised by three Euler angles: $\alpha_{q}$, $\beta_{q}$, and $\gamma_{q}$:
@@
In practice, when DM distributions are derived from observational data, the orientation angles and intrinsic shapes of DM are highly degenerate. To simplify this, we set $Z=z_{\rm gc}$ by assuming $\beta_{q}=0$. The orientation of the DM halo in the $x_{\rm gc}y_{\rm gc}$ plane is left unrestricted. By projecting DM particles along the $Z$ axis, we determine the principal axis in the $XY$ plane using particles within $r<100$ kpc. The long axis in this plane is labelled as $X$ and the short axis as $Y$. We define the axis ratio as $p_{\rm DM}=Y/X$, and $q_{\rm DM}=Z/X$. In this definition, $p_{\rm DM}$ is restricted to be <1<1, while $q_{\rm DM}$ is allowed to be >1>1.
The intrinsic shape of Auriga halo 23, 5, and 12 measured in such a coordinate $XYZ$ is shown in Fig. 1. The intrinsic shapes measured in our coordinate XYZXYZ are slightly different but close to the intrinsic shapes measured along the principal axes of the DM halo shown in Prada et al. (2019).
@@ 2.2 Mock data
By cross-matching LAMOST with Gaia DR3, we have a sample of 619,284 k-giants with 3D velocity measured, the typical errors of radial velocity $v_{r}$, tangential velocity $v_{\phi}$, $v_{\theta}$ are 10, 30, 30 km/s, respectively. The typical distance error is about $15\%$. We exclude disk stars chemodynamically to avoid biasing the spatial distribution of the halo. We carefully exclude all substructures, including streams, overdensities, and GES using the friend-of-friend method (Wang et al., 2022; Yang et al., 2019), and further perform $3\sigma$ clipping to remove outliers. In the end, we remain with 15,000 k-giants in the smooth halo with the spatial distribution shown in Fig. 2.
@@
First, we treat each particle as a single star, transferring their positions and velocities into the ICRS coordinate. We adopt a relative error of $15\%$ for distance, an error of 30 km/s for the tangential velocity in the RA and DEC directions, and an error of 20 km/s for the LOS velocity. The positions and velocities of stars are then perturbed by random numbers, normally distributed with dispersions equal to the observational errors.
Second, we perform spatial selection of the stars. We divide the particles into an original bulge with $r_{\rm gc}<4$ kpc, an original disk with $z_{\rm gc}<4$ kpc, $r_{\rm gc}>4$ kpc and $Z/Z_{\odot}>0.5$, and an original halo with the rest. We first randomly select 1/100 particles from the disk, 1/3 particles from the halo (1/4 for Auriga 12), and no particles from the bulge, then perform further spatial selection by only including stars in the northern hemisphere and within a spatial volume similar to the MW LAMOST + Gaia observations, as shown in Fig.2. This results in $\sim 30000$ stars with the full 6D phase-space information measured. The sample size is much smaller than the original LAMOST + Gaia sample because we take a sparse sampling of the disk stars.
@@ 2.3 Data preparation and correction of selection function
With mock data created in an observational manner, we convert the stars into the galactocentric coordinate and obtain the uncertainty of position and velocity in the new coordinate by a bootstrapping process. We clean the data and correct the select function to make it ready for the model construction, in a way similar to how we deal with the real observations.
@@
First, we use a position-velocity clustering estimator method to identify substructures; we select the smooth halo by removing all stars identified in substructures with a link length of 0.84 in the 6D distance (Yang et al., 2019). The stellar halos of Auriga 5 and 23 are rather smooth with no significant structures; we only perform this step on the mock data of Auriga 12. There are fewer substructures in the halo of Auriga simulations than in the real Milky Way galaxy due to the limited mass resolution in the simulation.
[…]
@@
Third, we clean the data by $3\sigma$ clipping in the velocity distributions. Although we have removed the substructures identified by the clustering estimator method, there are still some particles with high tangential velocities which are unlikely to be part of a smooth halo in dynamical equilibrium. We also performed a $3\sigma$ clipping in $v_{\phi}$ versus $v_{\theta}$ to remove the outliers.
@@
Lastly, we correct for the selection function. For this proof-of-concept, we take a simple approach by correcting the selection function in the $R_{\rm gc}-z_{\rm gc}$ plane. We divide the $R_{\rm gc}-z_{\rm gc}$ plane into $50\times 50$ bins in $R_{\rm gc}=[0,50]$ kpc and $z_{\rm gc}=[0,50]$ kpc. We calculate the density of stars in each volume of $2\pi R_{\rm gc}dR_{\rm gc}dz_{\rm gc}$ constructed from our sample ($\rho_{\rm select}$) and that of the original halo before performing spatial selection ($\rho_{\rm original}$). Any star $i$ in the bin $j$ will have a weight of $w_{i}=\rho_{\rm original,j}/\rho_{\rm select,j}$. Such a correction of selection is less accurate at $z_{\rm gc}<4$ kpc where disk stars are mixed in the original halo.
@@
In the end, we have about 20,000 stars in the smooth halo; these stars are all located in the northern hemisphere and have the selection function corrected. The final sample of stars with their weights is shown in Fig. 2.
@@
The real selection function of LAMOST is complicated and is corrected by comparing with photometric data which are considered complete with the magnitude limit (Liu et al., 2017; Yang et al., 2022), resulting in the correction of the selection function in the 3D space, more accurate than what we did here. It is beyond the scope of this paper to test the method of selection function correction. We just wanted the spatial sampling of our mock data to match some major properties of the real observations: (1) the data are not complete in different azimuthal angles, and we have less data in the direction of the Galactic Centre; (2) the observations at larger distances or with small polar angles are less complete; thus, we got higher weights for the stars at these positions. As shown in Fig. 2, the particle weight distribution in our mock data set generally matches the trend in the LAMOST + Gaia MW data.